Ten years ago Mike was one and a half times the age of Julia. Now Mike is 5 years older than Julia. What is Julia's age?
Answer by addingup(3677)

M-10= 1.5(J-10)
M= J+5
J+5-10= 1.5J-15
J-5= 1.5J-15
-.5J= -10
J= 20 Julia is 20 and Mike is 20+5= 25
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check - the problem says:
"Ten years ago Mike was one and a half times the age of Julia"
Mike: 25-10=15
Julia: 20-10=10
15/10= 1.5 We have the correct answer
J